Mighty Safe Fire Alarm is currently buying 59,000 motherboards from MotherBoard, Inc. at a price of $68 per board. Mighty Safe is considering making its own boards. The costs to make the board are as follows: direct materials, $32 per unit; direct labor, $12 per unit; and variable factory overhead, $14 per unit. Fixed costs for the plant would increase by $87,000. Which option should be selected and why?

a.buy, $87,000 more in profits

b.make, $503,270 increase in profits

c.make, $590,000 increase in profits

d.buy, $503,270 more in profits
Answer #1

Cost of buying = (59,000 * 68) = 4,012,000

Cost of making = (32+12+14)*59,000 + 87,000 = 3,509,000

So, it is better to make as it saves 503,000 in profits

Option B is the answer